A company imports a product from China at a cost of $50 per unit. The product is imported with an import tariff of 20% and it is sold for $100 per unit. The total sales revenues for the company were 85.000. If the government increases the import tariff per unit of the product to 30% and the elasticity of demand is 3 find the following:


(i) How many units of this product the company will sale if the price rises the same amount with the import tariff (20 Marks)


(ii) Calculate if the revenues of the company will increase or decrease with the increase of import tariff (20 Marks)



Explain each breakdown, please the total revenues that have 85.000 there is a decimal.


1
Expert's answer
2022-02-10T09:47:29-0500

"Price Elasticity=\frac{∆InQuantityDemanded}{∆InPrice}"

Making ∆ in Quantity the subject of the formula:

"\u2206In Quantity=Price Elasticity\u00d7 \u2206InPrice"

"=3\u00d7(30-20)=30" %

"Total UnitsSold= \frac{Total Revenue}{Price PerUnit}"

"=\frac{85000}{100}=850 Units"

Since the quantity demanded will decrease by 30% due to the price increase, the firm will sell:

"\frac{850×(100-30)}{100}= 595Units"

II) Since the price has increased by 10%, the new price will be:

"\frac{100×(100+10)}{100}=110" $

The new total revenue will be"=P\u00d7Q=110\u00d7595=65,450" $

Therefore the total revenue will decrease.